Solve:

a. the quadratic equation

b. the quadratic inequality


1
Expert's answer
2021-10-11T01:26:20-0400

For example:


a.

"6x^2-7x+2=0"

"x=\frac{7\pm \sqrt{49-48}}{12}"

"x_1=1/2,\ x_2=2/3"


b.

"6x^2-7x+2<0"

"6x^2-7x+2=6(x-1\/2)(x-2\/3)=(2x-1)(3x-2)"

"(2x-1)(3x-2)<0"

"x\isin (1/2,2/3)"